OpenStudy (anonymous):

39.75 into a mixed number

OpenStudy (anonymous):

\[39.75=39+\frac{75}{100}=\frac{159}{4} \]

OpenStudy (anonymous):

Use a calculator to verify that 159/4 = 39.75

OpenStudy (anonymous):

Thank you for the medal.